OpenStudy (anonymous):

x=sec squret-1,y=tant,t=-pi/4

hartnn (hartnn):

what do you need to do with that ?

hartnn (hartnn):

all i can think of is, \(x = \sec^2 t -1 = \tan^2 t = (\tan t)^2 = y^2 \)

hartnn (hartnn):

or do you need the derivative dy/dx at t= -pi/4 ??